# Higher-order homogenised riblet boundary conditions

### v1

@@ 7 Summing-up and numerical examples
An even larger number of terms are zero in the boundary conditions for the slip-no-slip surface, for the reason that some of the polynomial behaviours required at $Z\rightarrow\infty$ already have $u$ and its derivative $u_{z}$ (or $v$ and $v_{z}$) simultaneously vanish at the wall, and therefore satisfy the boundary conditions on both the slip and no-slip sections of the wall without any $Y$-dependence required. Such polynomial solutions do not entail any corrections and produce zero coefficients. In fact, as can be seen from figures 3 and 4, all the second-order velocity fields in the last column are either constant with $Y$ or of odd parity. Astoundingly enough, no second-order protrusion heights and only 8 third-order protrusion heights have non-zero values in the slip-no-slip case of table 1. Among these all coefficients involved in the $w(x,y,0,t)$ boundary condition (bottom row of (7.1)) are zero, for the trivial reason that the boundary condition $w(x,y,0,t)=0$ is exact.
@@ 8 Conclusions
As discussed in the introduction, the original concept of longitudinal and transverse protrusion heights, and the connection of their difference with the initial slope of the drag-reduction curve of riblets, have been extended by several authors over the years, both in order to further inform the design of drag-affecting surfaces by predicting their performance beyond the linear regime, and as equivalent slip-length boundary conditions to simplify numerical simulations of the same. This paper is an attempt to systematize such extensions by continuing the asymptotic expansion, of which the protrusion heights represent the first order, to higher orders, and in the while to discover the role of nonlinearities. The appropriate expansion parameter is unequivocally $\varepsilon=s^{+}$, the riblet period in wall units. The strength and at the same time the weakness of this approach is that it is a formal asymptotic expansion in the limit of $\varepsilon\rightarrow 0$, so on the positive side every term is strictly identified and there is no ambiguity as to what is negligible compared to what, on the negative side there is no guarantee that the approximation will be useful for practically interesting values of $\varepsilon$ (i.e., $s^{+}$). Nonetheless, there also is no guarantee for empirical approximations not based on a formal expansion, and at the very least the availability of approximations of multiple orders allows one to estimate a posteriori how close we are to reality by comparing successive orders.
@@
The main operational results, at a first reading that skips the mathematics involved, are collected in the third-order equivalent boundary condition (7.1) and the table of its coefficients for five typical wall textures, table 1. We recall that the purpose of such an equivalent boundary condition is to replace the actual fine wall geometry, in the small-roughness limit where all of its dimensions tend to zero simultaneously, by a virtual flat wall where an approximate boundary condition is imposed. The advantage for numerical simulations, which otherwise become increasingly difficult as the texture becomes smaller and smaller in size, should be obvious.
@@
Already deciding the appropriate, most general form of such an equivalent boundary condition was not evident before we started the present work. The practical answer is contained in the Cauchy expansion of an initial-value problem: a partial differential equation, or a system of those, can at least formally be specified as an initial-value problem in the wall-normal coordinate which for the Navier-Stokes equations involves six quantities, three components of the velocity vector and three components of the normal-stress vector, and their surface-parallel and time derivatives. The most general boundary condition that transforms this into an elliptic boundary-value problem involves three constraints among these six quantities. Most convenient is to write these three constraints in explicit form by assigning the three velocity components as functions of the three stress components and their wall-parallel derivatives, so that the solid-wall condition of zero velocity is naturally recovered when texture size $s^{+}$ tends to zero. This is what (7.1) does. One must note that, whereas wall-normal derivatives of stresses (equivalently replaced by $u_{z}$, $v_{z}$ and $p$) could in principle also appear in this formula, they can always be algebraically eliminated by using the differential equations themselves and their derivatives. Therefore a canonical form where wall-normal derivatives of $u_{z}$, $v_{z}$ and $p$ never appear can and should be used to eliminate ambiguities. This is notably also true of the wall-normal derivative $w_{z}$ of the wall-normal velocity, which can always be eliminated through the continuity equation and should not appear in the equivalent boundary condition at all.
@@
The analysis proceeded through several intermediate problems, starting with the 2D Laplace and the 2D Stokes problem (which, not by accident, were the basic ingredients of the 1991 protrusion-height theory as well), and discovering step by step that these are also the building blocks of both 2D and 3D Navier-Stokes problems, when the wall geometry itself is 2D (as it is in classical riblets but not, for instance, in recent attempts at wavy riblets, or in general roughness). The limitation of the present article to 2D geometries, which anyway include a large number of practically interesting cases, is intentional as it gives the theory interesting distinctive features, but is not insurmountable. Effective boundary conditions for more general 3D textures were already introduced at first order by Sarkar & Prosperetti (1996) and by Luchini (2013), in the small-roughness limit, and at second order in the shallow-roughness limit by Kamrin et al. (2010), and could be generalized to higher orders in the future.
@@
One outcome was to clarify, if there was any need, why only one longitudinal and one transverse protrusion heights were included in the (Luchini et al., 1991) original theory, despite the presence of two wall boundary conditions in the cross-flow problem. A formalization that specifies the most general relationship between the zeroth and first derivatives of the streamfunction $\psi$ on one side and its second and third derivatives on the other, through a matrix of four $a,b,c,d$ coefficients (4.3), shows that these coefficients are of different orders in the scale ratio $\varepsilon$, and $a$ alone has a role at first order. The $b,c,d$ coefficients only appear at higher orders, and were appropriately omitted by Luchini et al. (1991). They do play a role here, on the other hand.
@@
Among the qualitative conclusions that can be drawn of the present analysis, even more interesting than what is included in the effective boundary condition (7.1) probably is what is not there. In particular, symmetry considerations forbid about half of possible couplings when riblets are left-right symmetric, because these coefficients would mix incompatible symmetries if present. We already mentioned that normal derivatives different from the basic shear components $u_{z}$, $v_{z}$ have no role in it, not so much because they cannot be there but because they need not be there; the Cauchy expansion process can always replace them with other quantities. In addition, undifferentiated pressure never appears (not so surprisingly because in incompressible flow pressure is defined up to an additive arbitrary constant), and this is another reason why the first-order equivalent boundary condition, in the first block of (7.1), only contains two protrusion heights.
@@
At second order we see the appearance, still in the first line of (7.1), of an effect of the longitudinal and transverse pressure gradients on the respective wall-parallel velocity components, but also effects of shear-stress gradients $v_{yz}$ and $u_{xz}$ on the wall-normal velocity. All these effects have coefficients with the dimensions of an area, just because being of second order in the expansion they are proportional to the second power of the texture period $s^{2}$. The second derivative of wall normal velocity $w_{zz}$ need not appear because, from the continuity equation, it can always be replaced by $-u_{xz}-v_{yz}$. In addition, an unforeseen relationship was discovered between the effect of each pressure gradient on the corresponding velocity and the effect of the gradient of shear stress on normal velocity, in that their coefficients are necessarily equal, as evidenced in table 1 and justified in appendix A. The number of independent second-order protrusion heights for left-right symmetric riblets thus turns out to be just two. Two more second-order terms, only present for asymmetric riblets, involve cross effects of $v_{xz}$ upon longitudinal velocity $u$ and of $u_{xz}$ upon transverse velocity $v$, also with equal (aside from sign) coefficients. Two more coefficients, $h_{2}$ tying $u$ to $u_{yz}$, and $a_{2}$ tying $v$ to $v_{yz}$, will never be present for either symmetric or asymmetric riblets (see appendix B), despite these being the first that naturally turned up as “second-order protrusion heights” in the theory.
@@
Concerning the slip-no-slip striped surface that is the simplest idealized model of a superhydrophobic surface, we found an unexpected important simplification in that all second-order effects are zero and the first non-zero corrections to the standard protrusion heights are of the third order. With hindsight this lack of second-order effects can be traced back to the lack of possible actions upon the normal velocity, because the impermeability condition is satisfied exactly, combined with the property above that states that the effects of pressure gradients have the same weights as the effects upon normal velocity.
@@
But our perhaps most striking conclusion is that, amid the numerous third-order effects that are listed in table 1 and will not be repeated here, nonlinear terms are missing. Already that nonlinearities of the Navier-Stokes equations should not turn up before the third order was not obvious before the present formalization of the matched asymptotic expansion, but further when they do show up at third order they do so with a zero coefficient, not only for symmetric riblets (where these terms would be forbidden by symmetry), but also for any geometry as shown in appendix B. Therefore nonlinearity is repelled to fourth order (at least) in $s^{+}$, and linear equivalent boundary conditions can work unencumbered up to third order.
@@
Acknowledgements. P. Luchini wishes to thank the Department of Mechanical Engineering of the University of Melbourne for a two-month invitation (March-April 2024) during which the core of this work was conceived.
@@
Funding. D. Chung acknowledges support by the Air Force Office of Scientific Research (AFOSR) under award number FA2386-23-1-4071 (AOARD program manager: David Newell).

### v2

@@ 7 Summing-up and numerical examples
An even larger number of terms are zero in the boundary conditions for the slip-no-slip surface, for the reason that some of the polynomial behaviours required at $Z\rightarrow\infty$ already have $u$ and its derivative $u_{z}$ (or $v$ and $v_{z}$) simultaneously vanish at the wall, and therefore satisfy the boundary conditions on both the slip and no-slip sections of the wall without any $Y$-dependence required. Such polynomial solutions do not entail any corrections and produce zero coefficients. In fact, as can be seen from figures 4 and 5, all the second-order velocity fields in the last column are either constant with $Y$ or of odd parity. Astoundingly enough, no second-order protrusion coefficients and only 8 third-order protrusion coefficients have non-zero values in the slip-no-slip case of table 1. Among these, all coefficients involved in the $w(x,y,0,t)$ boundary condition (bottom row of (7.1)) are zero, for the trivial reason that the boundary condition $w(x,y,0,t)=0$ is exact.
[…]
homogenisation error
$\varepsilon$
Figure 9: Homogenisation error of the equivalent boundary conditions (7.1), calculated for sawtooth riblets with $U_{Z}(0)=0.480125-0.754587\mathrm{i},V_{Z}(0)=0.268092-1.27046\mathrm{i},P(0)=0.592799-0.367651\mathrm{i},\alpha=0.835223,\beta=0.777775,\omega=1.26823$ (see Appendix C for the precise definition of these quantities).
A word is also needed about the accuracy of formula (7.1) and of the coefficient values in Table 1, not just intended as numerical accuracy but as freedom from formulation or programming mistakes, since the formulation itself is complicated enough that to rely on one’s attention only to exclude inadvertent mistakes is insufficient. For this purpose a test case has been developed, consisting of a laminar flow past riblets with an imposed sinusoidal stress excitation. (Sinusoidal in all of the $x$, $y$, $t$ directions, and in all three surface stress components, so as to excite all possible degrees of freedom. For the same reason, of exercising all possible coefficients in (7.1), the test was conducted upon asymmetric sawtooth riblets. Needless to say, the error plot of symmetric riblets turns out similar.) A predefined sinusoidal stress vector was imposed on a plane at a certain distance (3 riblet periods) from the wall, and the resulting velocity vector on the same plane was measured. The 3D, unsteady Stokes problem for this flow was solved both exactly and in homogenised form, and the absolute value of the difference between the two resulting complex velocity vectors, $\left(\left|U_{\text{out,hom.}}-U_{\text{out,ex.}}\right|^{2}+\left|V_{\text{out,hom.}}-V_{\text{out,ex.}}\right|^{2}+\left|W_{\text{out,hom.}}-W_{\text{out,ex.}}\right|^{2}\right)^{1/2}$, was chosen to represent the homogenisation error. Nontrivial difficulties had to be overcome, related to the possible interference of discretization error with the error of the expansion we are trying to measure, the details of which are discussed in Appendix C. The results of this test are displayed in Figure 9.
[…]
Figure 9 also gives the interesting indication that the present expansion can be practically useful in a more and more extended range of values of $\varepsilon$ with higher and higher order, as it can be seen that the error curve for third order is consistently below the one for second order, and the error curve for second order is consistently below the one for first order, over the whole range of $\varepsilon$. This was not a priori granted: it is not uncommon for an asymptotic expansion to exhibit a shorter validity range with increasing order, but evidently the present case has favourable behaviour.
7.1 Comparison with previous literature
Up to the second order, the asymptotic expansion (7.1) is consistent with the two-dimensional $(y,z)$, two-component $(v,w)$ rough-wall equations (7) and (8) and table 5 of Sudhakar et al. (2021), written here in the present notation and wall scaling (but without changing the numerical subscripts of the coefficients):
$v=\varepsilon[\mathcal{L}_{11}(v_{z}+w_{y})]+\varepsilon^{2}[\mathcal{K}_{11}(-p_{y}+2w_{zy})],\quad w=\varepsilon^{2}[\mathcal{M}_{211}(v_{zy}+w_{yy})].$ (7.2)
Following Appendix A of Ahmed and Bottaro (2025), we note that $w=\mathrm{O}(\varepsilon^{2})$ to neglect the $y$-derivatives $w_{y}$, $w_{zy}$ and $w_{yy}$ on the right-hand sides, and then identify, by matching with (7.1) presently, that $\mathcal{L}_{11}=a_{1}$, $\mathcal{K}_{11}=b_{2}$, $\mathcal{M}_{211}=-c_{2}$. Noting that $a_{2}=0$ (Appendix B) and ignoring the third component $u$ or third dimension $x$ in (7.1), it can be seen that (7.2) and (7.1) are identical up to the second order.
Likewise, the present (7.1) is consistent with the two-dimensional $(y,z)$, two-component $(v,w)$ third-order effective boundary conditions in equations (54) and (55) of Bottaro and Naqvi (2020), written in the present notation and wall scaling (without changing the numerical subscripts of the coefficients and dropping the superscripts of the coefficients):
$v=\varepsilon[\lambda(v_{z}+w_{y})]+\varepsilon^{2}[m_{12}(-p_{y}+2w_{zy})]$
$\quad+\varepsilon^{3}[\theta(v_{zyy}+w_{yyy})+q_{12}(v_{zt}+w_{yt})],$
$w=\varepsilon^{2}[m_{21}(v_{zy}+w_{yy})]+\varepsilon^{3}[q_{21}(-p_{yy}+2w_{zyy})].$
First, up to second order, it can be verified that this is consistent with (7.2) above, i.e. the equation of Sudhakar et al. (2021), with $\lambda=\mathcal{L}_{11}$, $m_{12}=\mathcal{K}_{11}$ and $m_{21}=\mathcal{M}_{211}$ and, therefore, also consistent with (7.1). Further, it has been recognised, e.g. table 3 of Bottaro and Naqvi (2020), that $m_{21}=-m_{12}$, i.e. $\mathcal{K}_{11}=-\mathcal{M}_{211}$, consistent with the numerical values in table 5 of Sudhakar et al. (2021), and also, in the present notation, $b_{2}=c_{2}$, also verified in the present table 1 and now proved in Appendix A. Focussing now on the third-order terms, noticing again that $w=\mathrm{O}(\varepsilon^{2})$, such that the $y$-derivatives $w_{yyy}$, $w_{yt}$ and $w_{zyy}$ can be neglected, we then identify by comparing with (7.1) that $\theta=a_{3}$, $q_{12}=e_{3}^{(v_{zt})}$ and $q_{21}=d_{3}$ in addition to $\lambda=a_{1}$, $m_{12}=-b_{2}$ and $m_{21}=-c_{2}$ already established at second order. The remaining unmatched, extra terms in (7.1) are either zero for symmetric shapes (indicated in grey) or involve the third dimension $x$ or the third component $u$. The third-order coefficient $e_{3}^{(nl)}$ is zero (Appendix B), but the third-order coefficient $b_{3}=c_{3}$ (cf. Appendix A), multiplying $p_{yy}$ and $v_{yyz}$, is non-zero for asymmetric shapes (table 1), but not previously discussed, e.g. by Bottaro and Naqvi (2020). The claim that $q_{12}=q_{21}$ (Bottaro and Naqvi, 2020), i.e. that $e_{3}^{(v_{zt})}$ and $d_{3}$ are equal is not supported by the present data (table 1).
So far, no coefficients have been compared numerically with previous literature. We address this now through the three-dimensional second-order boundary condition for a rough wall, (2.21) of Ahmed and Bottaro (2025) written with the present notation and wall scaling:
$u=\varepsilon[\lambda_{x}(u_{z}+w_{x})]+\varepsilon^{2}[\mathcal{K}_{xz}^{itf}(-p_{x}+2w_{zx})],$
$v=\varepsilon[\lambda_{y}(v_{z}+w_{y})]+\varepsilon^{2}[\mathcal{K}_{yz}^{itf}(-p_{y}+2w_{zy})],$
$w=\varepsilon^{2}[-\mathcal{K}_{xz}^{itf}(u_{zx}+w_{xx})-\mathcal{K}_{yz}^{itf}(v_{zy}+w_{yy})],$
cf. (68) and (69) of Naqvi and Bottaro (2021). For a rough wall, $\mathcal{K}_{zz}=0$ (Ahmed and Bottaro, 2025, Appendix C). These 4 coefficients $(\lambda_{x},\lambda_{y},\mathcal{K}_{xz}^{itf},\mathcal{K}_{yz}^{itf})$ are to be compared with $(h_{1},a_{1},h_{2}^{(p_{x})}=g_{2},b_{2}=-c_{2})$ of (7.1), numerically tabulated presently in table 1 ($60^{\circ}$-triangular, rectangular, trapezoidal, sawtooth riblets) corresponding respectively to table 4 of Ahmed and Bottaro (2025) (equilateral triangle, thick blade, trapezoidal, right asymmetric triangle), showing that the numbers agree. At second order in three dimensions and three components, the new coefficient is $f_{2}=-h_{2}^{(v_{xz})}$, which is only non-zero for asymmetric shapes, multiplying $u_{xz}$ and $v_{xz}$, previously undocumented by either Ahmed and Bottaro (2025) or Sudhakar et al. (2021).
In summary, compared to previous literature, (7.1) highlights missing coefficients for asymmetric shapes. In particular, $f_{2}=-h_{2}^{(v_{xz})}$ was missing at second order for three-dimensional and three-component flow, and $b_{3}=c_{3}$ was missing at third order for two-dimensional and two-component flow. Moreover, a third-order boundary condition for three-dimensional and three-component flow had not been attempted, now addressed by (7.1), revealing an additional seven new coefficients $h_{3}$, $h_{3}^{(u_{zt})}$, $h_{3}^{(v_{xyz})}$, $h_{3}^{(u_{xxz})}$, $f_{3}^{(u_{xyz})}$, $f_{3}^{(v_{xxz})}$ and $g_{3}^{(p_{xx})}$ for symmetric shapes plus a further two, $h_{3}^{(p_{xy})}=-g_{3}^{(u_{xyz})}$ and $f_{3}^{(p_{xx})}=-g_{3}^{(v_{xxz})}$ (Appendix A) for asymmetric shapes. In (7.1), regardless of shape symmetry, $h_{2}=a_{2}=0$ and surprisingly $h_{3}^{(nl)}=e_{3}^{(nl)}=0$ (Appendix B), i.e. nonlinear terms are absent at third order.
@@ 8 Conclusions
As discussed in the introduction, the original concept of longitudinal and transverse protrusion heights, and the connection of their difference with the initial slope of the drag-reduction curve of riblets, have been extended by several authors over the years, both in order to further inform the design of drag-affecting surfaces by predicting their performance beyond the linear regime, and as equivalent slip-length boundary conditions to simplify numerical simulations of the same. This paper is an attempt to systematize such extensions by continuing the asymptotic expansion, of which the protrusion heights represent the first order, to higher orders, and in the while to discover the role of nonlinearities. The appropriate expansion parameter is unequivocally $\varepsilon=s^{+}$, the riblet period in wall units, as discussed by nondimensionalization arguments at the beginning of §5. The strength and at the same time the weakness of this approach is that it is a formal asymptotic expansion in the limit of $\varepsilon\rightarrow 0$, so on the positive side every term is strictly identified and there is no ambiguity as to what is negligible compared to what, on the negative side there is no guarantee that the approximation will be useful for practically interesting values of $\varepsilon$ (i.e., $s^{+}$). Nonetheless, there also is no guarantee for empirical approximations not based on a formal expansion, and at the very least the availability of approximations of multiple orders allows one to estimate a posteriori how close we are to reality by comparing successive orders.
@@
The main operational results, at a first reading that skips the mathematics involved, are collected in the third-order equivalent boundary condition (7.1) and the table of its coefficients for six typical wall textures, table 1. We recall that the purpose of such an equivalent boundary condition is to replace the actual fine wall geometry, in the small-roughness limit where all of its dimensions tend to zero simultaneously, by a virtual flat wall where an approximate boundary condition is imposed. The advantage for numerical simulations, which otherwise become increasingly difficult as the texture becomes smaller and smaller in size, should be obvious.
@@
Already deciding the appropriate, most general form of such an equivalent boundary condition was not evident before we started the present work. The practical answer is contained in the Cauchy expansion of an initial-value problem: a partial differential equation, or a system of those, can at least formally be specified as an initial-value problem in the wall-normal coordinate which for the Navier-Stokes equations involves six quantities, three components of the velocity vector and three components of the normal-stress vector, and their surface-parallel and time derivatives. The most general boundary condition that transforms this into an elliptic boundary-value problem involves three constraints among these six quantities. Most convenient is to write these three constraints in explicit form by assigning the three velocity components as functions of the three stress components and their wall-parallel derivatives, so that the solid-wall condition of zero velocity is naturally recovered when texture size $s^{+}$ tends to zero. This is what (7.1) does. One must note that, whereas wall-normal derivatives of stresses (equivalently replaced by $u_{z}$, $v_{z}$ and $p$) could in principle also appear in this formula, they can always be algebraically eliminated by using the differential equations themselves and their derivatives. Therefore a canonical form where wall-normal derivatives of $u_{z}$, $v_{z}$ and $p$ never appear can and should be used to eliminate redundance. This is notably also true of the wall-normal derivative $w_{z}$ of the wall-normal velocity, which can always be eliminated through the continuity equation and should not appear in the equivalent boundary condition at all.
The analysis proceeded through several intermediate problems, starting with the 2D Laplace and the 2D Stokes problem (which, not by accident, were the basic ingredients of the 1991 protrusion-height theory as well), and discovering step by step that these are also the building blocks of both 2D and 3D Navier-Stokes problems, when the wall geometry itself is 2D (as it is in classical riblets but not, for instance, in recent attempts at wavy riblets, or in general roughness). Our final result, the equivalent boundary condition (7.1) and the table of its coefficients, was confirmed by a numerical test case showing that the homogenisation error actually exhibits its theoretical convergence rate. Nontrivially, the homogenisation error of the higher-order condition is found to be lower not just asymptotically but in the whole range of $\varepsilon$ (and, implicitly, wavenumbers) examined, thus providing a uniformly better approximation.
@@
One may wonder how the present analysis of riblets, which are a very specific form of roughness, relates to more general 3D roughness. The limitation of the present article to spatially periodic 2D geometries, which anyway include a large number of practically interesting cases, is intentional as it gives the theory interesting distinctive features, but is not insurmountable. Effective boundary conditions for more general 3D textures were already introduced at first order by Sarkar and Prosperetti (1996) and by Luchini (2013), in the small-roughness limit, and at second order in the shallow-roughness limit by Kamrin et al. (2010), and could be generalized to higher orders in the future. We note, however, that the peculiarity of riblets, that the inner solution is 2D even when the outer flow is 3D, would be lost.
@@
One outcome was to clarify, if there was any need, why only one longitudinal and one transverse protrusion heights were included in the (Luchini et al., 1991) original theory, despite the presence of two wall boundary conditions in the cross-flow problem. A formalization that specifies the most general relationship between the zeroth and first derivatives of the streamfunction $\psi$ on one side and its second and third derivatives on the other, through a matrix of four $a,b,c,d$ coefficients (4.3), shows that these coefficients are of different orders in the scale ratio $\varepsilon$, and $a$ alone has a role at first order. The $b,c,d$ coefficients only appear at higher orders, and were appropriately omitted at first order by Luchini et al. (1991); they do play a role here, on the other hand.
@@
Among the qualitative conclusions that can be drawn of the present analysis, even more interesting than what is included in the effective boundary condition (7.1) is probably what is not there. In particular, symmetry considerations forbid about half of possible couplings when riblets are left-right symmetric, because these coefficients would mix incompatible symmetries if present. We already mentioned that normal derivatives different from the basic shear components $u_{z}$, $v_{z}$ have no role in it, not so much because they cannot be there but because they need not be there; the Cauchy expansion process can always replace them with other quantities. In addition, undifferentiated pressure never appears (not so surprisingly because in incompressible flow pressure is defined up to an additive arbitrary constant, and having undifferentiated pressure in the boundary condition would make this constant no longer arbitrary); this is another reason why the first-order equivalent boundary condition, in the first block of (7.1), only contains two protrusion heights.
@@
At second order we see the appearance, still in the first line of (7.1), of an effect of the longitudinal and transverse pressure gradients on the respective wall-parallel velocity components, but also effects of shear-stress gradients $v_{yz}$ and $u_{xz}$ on the wall-normal velocity. All these effects have coefficients with the dimensions of an area, just because being of second order in the expansion they are proportional to the second power $s^{2}$ of the texture period. The second derivative of wall normal velocity $w_{zz}$ need not appear because, from the continuity equation, it can always be replaced by $u_{xz}-v_{yz}$. In addition, an unforeseen relationship was discovered between the effect of each pressure gradient on the corresponding velocity and the effect of the gradient of shear stress on normal velocity, in that their coefficients are necessarily equal, as evidenced in table 1 and justified in appendix A. The number of independent second-order protrusion coefficients for left-right symmetric riblets thus turns out to be just two. Two more second-order terms, only present for asymmetric riblets, involve cross effects of $v_{xz}$ upon longitudinal velocity $u$ and of $u_{xz}$ upon transverse velocity $v$, also with equal (and opposite in sign) coefficients. Two more coefficients, $h_{2}$ tying $u$ to $u_{yz}$, and $a_{2}$ tying $v$ to $v_{yz}$, will never be present for either symmetric or asymmetric riblets (see appendix B), despite these being the first that naturally turned up as “second-order protrusion coefficients” in the theory.
@@
Concerning the slip-no-slip striped surface that is the simplest idealized model of a superhydrophobic surface, we found an unexpected important simplification in that all second-order effects are zero and the first non-zero corrections to the standard protrusion heights are of the third order. With hindsight this lack of second-order effects can be traced back for the normal velocity to its lack of any corrections because the impermeability condition is satisfied exactly, and for the pressure gradients to the above property which states that the second-order effects of pressure gradients have the same weights as the effects upon normal velocity.
@@
But our perhaps most striking conclusion is that, among the numerous third-order effects that are listed in table 1 and will not be repeated here, nonlinear terms play no role. Already that nonlinearities of the Navier-Stokes equations should not appear before the third order only becomes obvious after the present formalization of the matched asymptotic expansion; but even at third order, when nonlinearities do produce nontrivial inner velocity perturbations, their contributions to the equivalent boundary conditions vanish, for symmetric riblets because they are forbidden by symmetry, and more generally for any geometry as shown in appendix B. Therefore nonlinearity is repelled to fourth order (at least) in $s^{+}$, and linear equivalent boundary conditions can work unencumbered up to third order.
@@
[. Acknowledgements] P. Luchini wishes to thank the Department of Mechanical Engineering of the University of Melbourne for a two-month invitation (March-April 2024) during which the core of this work was conceived.
@@
[. Funding] D. Chung acknowledges support by the Air Force Office of Scientific Research (AFOSR) under award number FA2386-23-1-4071 (AOARD program manager: David Newell).